OpenStudy (anonymous):

help~ solving this for lambda1 and lamba2: 2+(lambda1)(2x)+(lambda2)=0 2+(lambda1)(2y)+(lambda2)=0 2z+(lambda1)(2z)+(lambda2)=0

OpenStudy (anonymous):

\(2+(\lambda_1)(2x)+(\lambda_2)=0\) \(2+(\lambda_1)(2y)+(\lambda_2)=0\) \(2z+(\lambda_1)(2z)+(\lambda_2)=0\) are you supposed to solve it in terms of x,y ,z?

OpenStudy (anonymous):

wow thank you I was just typing the equations out again. No the answer should be a number. and x+y+z=3

OpenStudy (anonymous):

and \[x ^{2} + y^{2} + z^{2} = 11\] I don't know if these have to be used...

OpenStudy (anonymous):

you're welcome :) oh gosh. lol these should help out a lot.

OpenStudy (anonymous):

does the third eq really have that 2z infront there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

may I ask.. can the lagrange multiplier be zero?

OpenStudy (anonymous):

they might? I'm not really sure. I've got (-1,1,3)for (x,y,z) if they are integers...but this isn't much help, is it?

OpenStudy (anonymous):

think I got it, thanks

OpenStudy (anonymous):

you're welcome :)